Arrow Crating & Packaging produces wooden packing boxes to be used in the military and industrial packaging industry. Current operations allow the firm to make 500 boxes per day, in two 8-hour shifts (250 boxes per shift). The firm has introduced some moderate changes in equipment, and conducted appropriate job training, so that production levels have risen to 300 boxes per shift. Labor costs average $10 per hour for each of the 5 full-time workers on each shift. Capital costs were previously $3,000 per day and rose to $3,200 per day with the equipment modifications. Energy costs were unchanged by the modifications, at $400 per day.

a) What is the firm's multifactor productivity before and after the changes?
Answer #1

Before the change,
daily Labour cost = no of workers*working hours per day*hourly wage = 5*(8*2)*10 = 800

Value of input = daily Labor cost + capital cost + energy cost = 800+3000+400 = 4200

Value of output = 500

Multifactor productivity = Value of output/Value of input = 500/4200 = 0.119047619 = 0.12 (Rounding to 2 decimal places)

After the change

daily Labour cost = no of workers*working hours per day*hourly wage = 5*(8*2)*10 = 800

Value of input = daily Labor cost + capital cost + energy cost = 800+3200+400 = 4400

Value of output = 500+300 = 800

Multifactor productivity = Value of output/Value of input = 800/4400 = 0.181818182 = 0.18 (Rounding to 2 decimal places)